OpenStudy (anonymous):

a ramp is extended from a truck to the ground. the ramp has a slope of 7/16. if y is 3.5 feet, what is , in feet?

OpenStudy (jonnyvonny):

We know the slope, m, is (7/16), and the y-intercept, b, is 0, for it starts from the ground. y=mx+b y=(7/16)x+0 y=(7/16)x We need to know what x value will obtain y=3.5, so get x alone. 3.5=(7/16)x What did you get?